OpenStudy (anonymous):

Convert 3 pi into radians?

OpenStudy (anonymous):

540° 500° 440° 400°

OpenStudy (zzr0ck3r):

\(3\pi\) is in radians. So you mean convert to degrees. Hint:\( \ \ \pi\rightarrow 180\)

OpenStudy (anonymous):

Yeah, that's why I didn't understand to question. Since it said convert to radians, but the answers were in degrees. thanks

OpenStudy (anonymous):

the question**

OpenStudy (anonymous):

It would be 540 correct?

OpenStudy (zzr0ck3r):

\(\pi \rightarrow 180\\2\pi \rightarrow 360\\3\pi \rightarrow ?\)

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

Thank you!